3(6x-1)=60
10-(x-2)=15
12-(x+3)=3
6-(3x-1)=19
9-(3x-2)=2
7-(3x-5)=13
5-(4x+3)=4
5x=3(x-6)
5x=2(x+12)
7(2x-3)=20
3(4x-2)=15x

Answer by BurningFlame_marc(77)

Hello!
I think you want me to distribute this for you. Okay I'll do it.
3(6x-1)=60---------18x-3=60
10-(x-2)=15--------10-x+2=15
12-(x+3)=3---------12-x-3=3
6-(3x-1)=19--------6-3x+1=19
9-(3x-2)=2---------9-3x+2=2
7-(3x-5)=13--------7-3x+5=13
5-(4x+3)=4---------5-4x-3=4
5x=3(x-6)----------5x=3x-18
5x=2(x+12)---------5x=2x+24
7(2x-3)=20---------14x-21=20
3(4x-2)=15x--------12x-6=15x

If you have any question, it will be my pleasure to answer it!
I hope you learned from me!